A ?typical? astronaut has a weight of 35 pounds-force on the moon. What is the weight of the astronaut on Mars, in units of pounds-force? The gravity on Mars is 3.7 meters per second squared.
Answer #1

here

the weight of astronaut on moon is

Fg = 35 pounds force = 155.68 kg m/s^2

gravity on moon is = 1.6 m/s^2

so by using the formula

Fmoon = m* g

155.68 = m * 1.6

m = 97.3 kg

so the mass of the astronaut is 97.3 kg

so then by using the formula

Fmars = m * a

Fmars = 97.3 * 3.7

Fmars = 360.01 kg m/s^2

Fmars = 80.933 lbf

so the weight on the mars is 80.933 lbf